OpenStudy (princesssleelee):

What is the exact value of x in the exponential equation 2^8x − 3 = 7? x = log4/8log2 x = 1/8log2 x= 5/8 x =log5/8

OpenStudy (wolf1728):

Tke logarithms of both sides

OpenStudy (wolf1728):

Take

OpenStudy (princesssleelee):

Im not exactly sure on how to do that

OpenStudy (wolf1728):

2^8x − 3 = 7 Okay, I'll take logarithms of both sides 2^8x = 10 log (2^8*x) = log(10) 8x * log(2) = log(10) 8x * 0.30102999566 = 1 Can you solve from here?

OpenStudy (wolf1728):

Maybe it should stay as 8x * log(2) = 1

OpenStudy (princesssleelee):

find x by solving: 8x * 0.30102999566=1?

OpenStudy (wolf1728):

OR 8x * log(2) = 1

OpenStudy (wolf1728):

Divide both sides by 8 * log(2)

OpenStudy (princesssleelee):

I solved that and got B :) is that correct?

OpenStudy (wolf1728):

Yes. See it's just that simple LOL

OpenStudy (princesssleelee):

the decimal really threw me off! lol

OpenStudy (wolf1728):

okkay - you've got it

OpenStudy (princesssleelee):

Do you mind helping me with a few more? :)

OpenStudy (wolf1728):

I think I see another problem that I could help someone else Maybe post the problem and I might get back

OpenStudy (princesssleelee):

Of course! Thank you! I will give you a medal each time!